OpenStudy (anonymous):

Tickets to the movies cost $12 for adults and $8 for children. A group of 14 people went to the movies, and the tickets cost $136. The system of equations models this situation, where x is the number of adults and y is the number of children. How many adults and how many children were in the group? A. 9 adults and 5 children B. 6 adults and 8 children C. 4 adults and 10 children D. 8 adults and 6 children

OpenStudy (anonymous):

the equation for this problem is: X+Y=14 12X +8Y=136

OpenStudy (anonymous):

can someone plz help me

OpenStudy (wilsondanielle):

this is more of a math question. However : If X is the number of adults and Y is the number of children, you know there has to be a total of 14 (equation #1) You also know that 12X + 8Y is equal to 136 (equation #2) because that is the amount that was send. The first step would be to isolate for one of the variables. From equation 1, X+Y=14. Therefore, X = 14-Y (and Y=14-X) Plug in either variable into equation #2 and solve for the remaining variable. \[12x+8(14-x)=136\] \[12x + 8*14 - 8x =136\] \[12x +112 - 8x = 136\] \[(12-8)x +112 = 136\] \[4x = 136-112\] \[4x = 24\] \[x = 24/4 = 6\] Therefore, 6 adults and 8 (14-6) children attended the movie.

TheSmartOne (thesmartone):

Please post this in the math subject. Thanks! http://www.openstudy.com/study#/groups/Mathematics